Calculate the number of electrons in 100grams  of co2
 

Grade:12

2 Answers

Arun
5 years ago
6 ele of C + 16 of O2 = 22 ele per molecule of CO2 

6.022 x 10²³ molecule = 44g mol⁻¹

moles of CO2 = 100 / 44

= 2.27 mol 

electron = 22 x 2.27 x 6.022 x 10²³ 
electron will be = 3x 10²⁵
Khimraj
5 years ago
electron in one mole of CO2 = 22*6.023*1023
So electrons in 100 gm or 100/44 mole of CO2 = (100/44)22*6.023*1023 = 3*1025
Hope it clears.